# Voost: A Unified and Scalable Diffusion Transformer for Bidirectional Virtual Try-On and Try-Off

### v1

@@ 4.3 Quantitative results
Methods VITON-HD [13] DressCode [55]
@@
StableVITON [41] 0.867 0.084 6.851 1.255 9.591 1.451 0.905 0.107 4.482 1.530 6.728 1.742
OOTDiffusion [84] 0.851 0.096 6.520 0.896 9.672 1.206 0.898 0.073 3.953 0.720 6.704 1.863
IDM-VTON [14] 0.881 0.079 6.343 1.322 9.613 1.639 0.923 0.048 3.801 1.201 5.621 1.554
CatVTON [15] 0.869 0.097 6.141 0.964 9.141 1.267 0.901 0.071 3.283 0.670 5.424 1.549
Leffa [91] 0.872 0.081 6.310 1.208 9.442 1.249 0.911 0.060 3.651 0.709 5.462 1.528
@@
Table 1: Quantitative results on VITON-HD [13] and DressCode [55] for the try-on task. We report both paired and unpaired evaluation results across benchmarks. Our unified dual-task model (Voost) consistently outperforms all baselines, including the single-task (VTON-only) model. Bold and underline indicate the best and second-best scores, respectively.
We evaluate visual fidelity and structural consistency using standard metrics. For realism, we report Fréchet Inception Distance (FID) [31] and Kernel Inception Distance (KID) [3]. To assess structural consistency, we use LPIPS [89] and SSIM [81]. As shown in Table. 1 and Table. 2, our model outperforms existing methods across all metrics.
@@ 4.4 User study
Figure 8: Qualitative comparison of try-on results with other commercial models  [23, 58, 43, 57, 61]. Best viewed in color and under zoom.
@@ 4.5 Ablation study
TryOffDiff [78] TryOffAnyOne [82] Ours (VTOFF-only) Ours
@@
Figure 10: Comparison results on in-the-wild images (Row 1–3) for the virtual try-off task. Unlike existing methods [78, 82], our method supports diverse garment types including upper, lower, and dresses (Last row).
@@
We compare training strategies with different subsets of trainable weights. As shown in Table 3, our attention-only tuning achieves the best overall performance, outperforming full fine-tuning, single-block training, and LoRA [33]. It effectively captures garment–person correspondence while significantly reducing training cost. Fig. 12 further highlights the qualitative advantage: attention-only tuning yields sharper, more coherent outputs with fewer artifacts. The supplementary material provides a detailed analysis of this design choice, highlighting sensitivity studies that justify our focus on attention layers.
@@ 5 Conclusion
Table 3: Quantitative comparison of training strategies with different updated layers on VITON-HD [13].
@@
In future work, we plan to incorporate additional cues such as body measurements or garment metadata to improve controllability. More broadly, the strong image-level foundation of Voost makes it well-suited for downstream extensions such as video-based [37, 29] or 3D-based [7, 30, 88] synthesis, where consistent and faithful garment–person interaction remains essential yet challenging.
@@ 6.1 Dataset composition
We train our model on a combination of VITON-HD [13], DressCode [55], and an additional curated set of real-world images. The latter consists of (1) high-resolution photographs captured in-house under controlled conditions and (2) in-the-wild fashion images from online sources. All external data were manually filtered to ensure high garment visibility, realistic lighting, and diverse poses.
@@ 6.2 Mask augmentation
For try-on, we ensure that all garment regions in the person image are fully masked, while explicitly preserving facial and hand regions. To obtain these masks, we leverage state-of-the-art human parsing and pose estimation models [42, 46, 25, 40]. During training, we generate diverse masks of varying shapes and sizes that perfectly occlude the clothing area but retain key human features. This encourages the model to rely entirely on the conditioning garment image rather than residual cues in the person input, significantly improving silhouette realism and garment fit, especially for longer garments.
For try-off, we use a dichotomous segmentation model [54] to isolate the garment region, and apply masks of varying sizes—from tight garment-specific masks to nearly full-image masks. This approach forces the model to infer plausible underlying appearances across a range of masking levels, enhancing robustness under challenging conditions.
@@ 6.3 Optimization
We use the AdamW [51] optimizer with a learning rate of $1\times 10^{-5}$ and a weight decay of 0.001. The model is trained with a batch size of 128 using the DeepSpeed ZeRO-2 [65] optimizer state partitioning strategy for approximately 48 single-H100 GPU days.
@@ 6.4 Inference
At test time, we employ a flow-matching-based Euler scheduler introduced in Stable Diffusion 3 [18], using exactly 28 sampling steps. All qualitative results shown in the supplementary figures were generated using this consistent inference setup.
@@ 6.5 Inference Technique Details
For self-corrective sampling, we apply updates at two timesteps, $t=5$ and $t=17$, based on a total denoising schedule of $T=28$. These correspond to early and mid-to-late stages of the denoising process, chosen to target complementary aspects: $t=5$ primarily influences coarse shape, while $t=17$ contributes to fine detail refinement. At the shape correction step ($t=5$), we use the same mask as the standard try-off setting, $\mathbf{M}=[\mathbf{1}\mid\mathbf{0}]$, encouraging the model to recover overall garment length and structure from the try-on result. For detail correction ($t=17$), we instead use a garment-specific mask obtained via segmentation models [54, 42, 40] of the input garment image. This focuses the gradient signal on texture and detail consistency, while avoiding penalization for plausible geometric differences. We set the number of refinement iterations to $R=5$ for each correction step.
@@ 7 Rationale for Attention-Only Training
Following [47, 44], we analyze parameter sensitivity by fully fine-tuning the model on concatenated inputs $\mathbf{X}_{\mathrm{g}}\mid\mathbf{X}_{\mathrm{p}}]$ and computing the normalized weight update for each layer as
@@ 11 More Qualitative Comparisons
We present additional qualitative comparisons with state-of-the-art baselines on VITON-HD [13], DressCode [55], and a set of in-the-wild examples gathered from online sources.

### v2

@@ 4.3 Quantitative results
Methods VITON-HD [choi2021viton] DressCode [morelli2022dress]
@@
StableVITON [kim2024stableviton] 0.867 0.084 6.851 1.255 9.591 1.451 0.905 0.107 4.482 1.530 6.728 1.742
OOTDiffusion [xu2024ootdiffusion] 0.851 0.096 6.520 0.896 9.672 1.206 0.898 0.073 3.953 0.720 6.704 1.863
IDM-VTON [idm-vton] 0.881 0.079 6.343 1.322 9.613 1.639 0.923 0.048 3.801 1.201 5.621 1.554
CatVTON [chong2024catvtonconcatenationneedvirtual] 0.869 0.097 6.141 0.964 9.141 1.267 0.901 0.071 3.283 0.670 5.424 1.549
Leffa [leffa] 0.872 0.081 6.310 1.208 9.442 1.249 0.911 0.060 3.651 0.709 5.462 1.528
@@
Table 1: Quantitative results on VITON-HD [choi2021viton] and DressCode [morelli2022dress] for the try-on task. We report both paired and unpaired evaluation results across benchmarks. Our unified dual-task model (Voost) consistently outperforms all baselines, including the single-task (VTON-only) model. Bold and underline indicate the best and second-best scores, respectively.
We evaluate visual fidelity and structural consistency using standard metrics. For realism, we report Fréchet Inception Distance (FID) [heusel2017gans] and Kernel Inception Distance (KID) [binkowski2018demystifying]. To assess structural consistency, we use LPIPS [zhang2018unreasonableeffectivenessdeepfeatures] and SSIM [wang2004image]. As shown in Table. 1 and Table. 2, our model outperforms existing methods across all metrics.
@@ 4.4 User study
Figure 8: Qualitative comparison of try-on results with other commercial models  [gemini2, chatgpt4o, kolors, vella, piccopilot]. Best viewed in color and under zoom.
@@ 4.5 Ablation study
TryOffDiff [tryoffdiff] TryOffAnyOne [tryoffanyone] Ours (VTOFF-only) Ours
@@
Figure 10: Comparison results on in-the-wild images (Row 1–3) for the virtual try-off task. Unlike existing methods [tryoffdiff, tryoffanyone], our method supports diverse garment types including upper, lower, and dresses (Last row).
@@
We compare training strategies with different subsets of trainable weights. As shown in Table 3, our attention-only tuning achieves the best overall performance, outperforming full fine-tuning, single-block training, and LoRA [hu2022lora]. It effectively captures garment–person correspondence while significantly reducing training cost. Fig. 12 further highlights the qualitative advantage: attention-only tuning yields sharper, more coherent outputs with fewer artifacts. The supplementary material provides a detailed analysis of this design choice, highlighting sensitivity studies that justify our focus on attention layers.
@@ 5 Conclusion
Table 3: Quantitative comparison of training strategies with different updated layers on VITON-HD [choi2021viton].
@@
In future work, we plan to incorporate additional cues such as body measurements or garment metadata to improve controllability. More broadly, the strong image-level foundation of Voost makes it well-suited for downstream extensions such as video-based [karras2024fashion, he2024wildvidfit] or 3D-based [cao2024gs, he2025vton, zhang2025robust] synthesis, where consistent and faithful garment–person interaction remains essential yet challenging.
[…]
@@ 6.1 Dataset composition
We train our model on a combination of VITON-HD [choi2021viton], DressCode [morelli2022dress], and an additional curated set of real-world images. The latter consists of (1) high-resolution photographs captured in-house under controlled conditions and (2) in-the-wild fashion images from online sources. All external data were manually filtered to ensure high garment visibility, realistic lighting, and diverse poses.
@@ 6.2 Mask augmentation
For try-on, we ensure that all garment regions in the person image are fully masked, while explicitly preserving facial and hand regions. To obtain these masks, we leverage state-of-the-art human parsing and pose estimation models [kirillov2023segany, li2019selfcorrectionhumanparsing, guler2018densepose, khirodkar2024sapiens]. During training, we generate diverse masks of varying shapes and sizes that perfectly occlude the clothing area but retain key human features. This encourages the model to rely entirely on the conditioning garment image rather than residual cues in the person input, significantly improving silhouette realism and garment fit, especially for longer garments.
For try-off, we use a dichotomous segmentation model [meyer2025benusingconfidenceguidedmatting] to isolate the garment region, and apply masks of varying sizes—from tight garment-specific masks to nearly full-image masks. This approach forces the model to infer plausible underlying appearances across a range of masking levels, enhancing robustness under challenging conditions.
@@ 6.3 Optimization
We use the AdamW [loshchilov2018decoupled] optimizer with a learning rate of $1\times 10^{-5}$ and a weight decay of 0.001. The model is trained with a batch size of 128 using the DeepSpeed ZeRO-2 [rasley2020deepspeed] optimizer state partitioning strategy for approximately 48 single-H100 GPU days.
@@ 6.4 Inference
At test time, we employ a flow-matching-based Euler scheduler introduced in Stable Diffusion 3 [esser2024scalingrectifiedflowtransformers], using exactly 28 sampling steps. All qualitative results shown in the supplementary figures were generated using this consistent inference setup.
@@ 6.5 Inference Technique Details
For self-corrective sampling, we apply updates at two timesteps, $t=5$ and $t=17$, based on a total denoising schedule of $T=28$. These correspond to early and mid-to-late stages of the denoising process, chosen to target complementary aspects: $t=5$ primarily influences coarse shape, while $t=17$ contributes to fine detail refinement. At the shape correction step ($t=5$), we use the same mask as the standard try-off setting, $\mathbf{M}=[\mathbf{1}\mid\mathbf{0}]$, encouraging the model to recover overall garment length and structure from the try-on result. For detail correction ($t=17$), we instead use a garment-specific mask obtained via segmentation models [meyer2025benusingconfidenceguidedmatting, kirillov2023segany, khirodkar2024sapiens] of the input garment image. This focuses the gradient signal on texture and detail consistency, while avoiding penalization for plausible geometric differences. We set the number of refinement iterations to $R=5$ for each correction step.
@@ 7 Rationale for Attention-Only Training
Following [li2020fewshotimagegenerationelastic, kumari2023multiconceptcustomizationtexttoimagediffusion], we analyze parameter sensitivity by fully fine-tuning the model on concatenated inputs $\mathbf{X}_{\mathrm{g}}\mid\mathbf{X}_{\mathrm{p}}]$ and computing the normalized weight update for each layer as
@@ 11 More Qualitative Comparisons
We present additional qualitative comparisons with state-of-the-art baselines on VITON-HD [choi2021viton], DressCode [morelli2022dress], and a set of in-the-wild examples gathered from online sources.